x^2 - 3x - 28
----------------
x^2 - 9x + 14 Assume that the denominator is not equal to zero.

Answer by richwmiller(9144)

Try to find factors for both num. and den. and cancel out the same.
(x-7)*(x+4)=x^2-3x-28
(x-7)*(x-2)=x^2-9x+14
so we end up with
(x+4)/(x-2)